OpenStudy (anonymous):

Using the discriminant, what values of c will make the quadratic y = -4x2 + 5x + c have no real solutions? A) (-∞, -25/16) B) (-∞, -25/16] C) (-25/16, ∞) D) [-25/16, ∞)

zepdrix (zepdrix):

\(\large y=-4x^2+5x+c\) Using the Quadratic Formula: \(\large x=\dfrac{-b\pm\color{orangered}{\sqrt{b^2-4ac}}}{2a}\) We only care about this orange part, the discriminant. Plugging in our values gives us,\[\large \sqrt{b^2-4(a)(c)}\qquad \rightarrow\qquad \sqrt{5^2-4(-4)(c)} \qquad \rightarrow\qquad \sqrt{25+16c}\] So what they're asking essentially is, `what values of c will cause us to take the square root of a negative number`?